# EnTruth: Enhancing the Traceability of Unauthorized Dataset Usage in Text-to-image Diffusion Models with Minimal and Robust Alterations

### v1

@@ 1 Introduction
The latest advancements in generative diffusion models (GDMs) [1, 2, 3], especially the text-to-image (T2I) models [4, 5] which excel in creating high-quality images that closely align with the given textual prompts, have revolutionized the field of image generation. These advantages stem not only from the development of model architectures and computing power, but also from the availability of large-scale datasets [6, 7, 8]. While datasets play an important role, their copyright protection has remained as an unsolved issue. The protection of these datasets’ copyrights is paramount for multiple reasons. For instance, open-source datasets [9] are generally available only for educational and research purposes, barring any commercial use. Additionally, for commercial datasets, it is crucial for companies to secure them from theft and unauthorized sales. While pre-training and fine-tuning both raise concerns of copyright infringement, fine-tuning has a more severe impact on the copyright of datasets. Compared to pre-training, fine-tuning is highly efficient, allowing for many unauthorized uses without effective regulatory restrictions.
Observing the above, techniques like watermarking [10, 11, 12, 13] and black-box Membership Inference (MI) [14, 15] have been employed to protect data specifically against unauthorized fine-tuning in text-to-image diffusion models. Nevertheless, existing watermark methods often face some common problems. For example, they usually modify a large portion [12] or even the whole of the dataset [11], which is not realistic for large-scale datasets. They also unexpectedly affect the quality of generation and are not robust enough under image corruption [13, 11]. Meanwhile, as black-box MI does not alter the data to boost the detection, it needs highly extensive queries to get a significant result. Another line of techniques, poison-only backdoor attack [16, 17], can be adapted for detecting dataset usage by verifying the attacked behavior. However, they are inherently designed for malicious attacking and demonstrate reduced robustness when subjected to re-captioning (as shown by Sec 5.2).
@@
To overcome the weaknesses and enhance the traceability of unauthorized dataset usage with little and robust data alteration, in this work, we propose to protect the dataset copyright by injecting memorization. In T2I models, memorization refers to the phenomenon where the models memorize and reproduce training examples when queried by a memorized prompt [18, 19, 20]. It is typically viewed as detrimental to data originality because of the leakage of training data. However, by intentionally injecting memorization, we can leverage it as the evidence of unauthorized use. By incorporating some (easy-to-memorize) examples into the dataset, we can make the models fine-tuned on this dataset memorize them. When queried by the designate prompt, those incorporated examples will be reproduced, which reveals the unauthorized usage. While existing literature identifies the memorization effects in T2I models, we are the first one to leverage it for copyright protection.
@@
According to whether the training examples are partially or entirely memorized, memorization can be divided into exact memorization (EM) and template memorization (TM) [21, 22]. To compare EM and TM, EM is the easier one to inject since it is found that simple duplicate data can cause EM [18, 23]. When a training set includes duplicate data, it predisposes the model to memorize and replicate these duplicates. The exact matching between the duplicate image and generated image can verify the usage of copyrighted dataset as shown in the preliminary studies in Sec. 3. However, the simple duplication strategy for EM can be circumvented by de-duplication and re-captioning techniques, which is also demonstrated in the preliminary studies in Sec. 3. In terms of TM, as shown in Fig. 1, the memorized training images share a common region (named as template), while their remaining areas (named as foreground) differ. Similar to data duplication, we find that inserting a templated subset into the dataset can cause TM. Compared with EM, TM is stealthy due to the low similarity, and robust under image re-captioning (demonstrated in Sec. 4.2 and Sec. 5.2).
@@
Observing the above difference between EM and TM, to generate a stealthy and effective templated set, we propose a novel framework, EnTruth, which Enhances the Traceability of unauthorized dataset usage by TM. Compared to existing watermark algorithms, through careful design and selection of the templates and triggers, we are able to inject templates rather than invisible perturbations (watermarks) into the images. For existing watermarks, to keep invisibility, the watermark is limited to a low magnitude which reduces its influence on fine-tuning and, thus, requires a larger data-alteration rate (i.e. modifying more data samples) as compensation. Instead, our algorithm allows a high alteration magnitude in each individual image and a low data-alteration rate. With such a design, we also enjoy two benefits. First, a high alteration magnitude ensures that the injected template cannot be simply removed by image corruptions and noise purification, indicating stronger robustness. Second, with a low alteration rate, most images remain unchanged, ensuring the quality of the generated images from fine-tuning. In addition to these key advantages, we accelerate memorization speed by controlling the similarity between the foregrounds of different images, strengthen robustness using soft triggers and further improve the watermark performance by multiple-query test. With EnTruth, the dataset owners can generate a templated set with a unique template and trigger token for their own dataset, which provides copyright protection with a low alternation rate, high accuracy, and robustness, without sacrificing the quality of generated images.
2 Related Works
Watermarks. Watermarking [24, 13, 11, 12, 10] is a widely used technique for tracing unauthorized data usage in diffusion models. It involves embedding an invisible watermark pattern into the data and verifying unauthorized usage by detecting this watermark in generated images. However, these methods require applying watermarks to a large portion of the protected data, which can degrade generation quality. Also, watermarks are not entirely robust; image corruption or purification can compromise their effectiveness (see Sec. 5.2).
Membership Inference. Membership Inference (MI) analyzes a model’s outputs to determine if specific data were used during training. MI can be categorized into white-box [25] and black-box [26, 15, 27, 14] settings. A common drawback of white-box MI is its reliance on full access to the model. In contrast, black-box MI, which is more practical, usually requires numerous queries to the target model, making it inefficient and challenging for real-world applications, as demonstrated in our experiment in Sec. 5.1.
Poison-only backdoor. Poison-only backdoor is designed to embed a detrimental behavior into a released model [28, 29, 30]. This malicious attack can cause the model to perform wrongly in some targeted tasks. For poison-only attacks [16, 17], it can be adapted to dataset protection by verifying the specific behavior. Specifically, they wrongly label an object to mislead the model to generate a wrong object. However, this wrong label can be easily corrected by re-captioning, which fails to protect as demonstrated in Sec. 5.2.
@@ 3 Preliminary Study
As mentioned in Section 1, memorization is a common phenomenon in GDMs, and we propose to leverage it in dataset protection. Depending on whether the generative images are totally or partially matching with the training images, memorization can be categorized into exact memorization (EM) and template memorization (TM), and the causes of them are different [22]. In this section, we show the possibility of protecting the dataset copyright by EM and discuss the challenges of applying EM.
@@ 3.1 Exact memorization by data duplication enhances the detection of unauthorized usage
Data duplication has been found as one important cause for exact memorization [18, 20]. By duplicating a specific data sample in the training set, the model can accurately memorize and generate it [23, 22]. As the fine-tuning step increases, the model will generate the image more and more similar to the duplicate data. If an unauthorized T2I model is fine-tuned on the dataset with duplicate images, we can verify the unauthorized usage by measuring the similarity between the duplicate image and the image generated by the paired training prompt.
In Fig. 2(a), we demonstrate the change of similarity score (measured by SSCD [31]) of duplicate data. We fine-tune Stable Diffusion (SD) starting from the checkpoint v1.4 using CC-20k, a subset of 20,000 text-image pairs from Conceptual Captions [7]. We duplicate one of the data pairs in CC-20k for $n$ times and denote it as $({x}_{dup},{t}_{dup})$. Usually, a larger $n$ can cause memorization with fewer steps. In Fig. 2(a), we use $n=32$. We denote other non-duplicate data as $({x},{t})$. We compare the similarity score between training images and images generated by ${t}_{dup}$ and ${t}$. In Fig. 2(a), the similarity score of duplicate data increases much faster than non-duplicate data. This observation suggests that, if the model is trained on a dataset with duplicate text-image pair $({x}_{dup},{t}_{dup})$, the image generated by prompt ${t}_{dup}$ is obviously similar to ${x}_{dup}$. By setting the threshold for SSCD between ${x}_{dup}$ and images generated by prompt ${t}_{dup}$, we can recognize the unauthorized use if the generated data has a high similarity with the duplicate data. Consequently, EM can achieve an accuracy of 74.5% at 10,000 fine-tuning steps with a threshold of 0.1 and 100% at 20,000 steps with a threshold of 0.2.
@@ 3.2 Challenges of Data Duplication
Data de-duplication. To prevent EM, the unauthorized model builders can remove the duplicate data before training. For example, Somepall et al. [20] calculate the similarity score, SSCD [31], of each pair of training images, and remove the cluster connected by high similarity scores. In Fig. 2(b), we plot SSCD of natural non-duplicate images. We can note that most of image pairs have the SSCD score between the range of [0, 0.2], while the duplicate data samples have the SSCD of 1. By setting a threshold of 0.7, which is a threshold commonly used to recognize identical images [20, 21, 23], all the duplicate data can be easily removed and no EM can be detected in generated images. Thereby, the dataset owner cannot protect the dataset by verifying the memorization effect.
Image re-captioning. EM relies on the memorized prompts to trigger the memorization. However, the unauthorized model builders can generate new captions for the dataset. Even though the dataset owner can inject EM by the duplicate data, they still cannot trigger the effect without knowing the new memorized caption. We generate new captions for cc-20k by BLIP [32], and fine-tune SD using the original dataset and the re-captioned dataset, respectively. In Fig. 2(c), we calculate SSCD between generated images and ${x}_{dup}$. When queried by original duplicate prompts (which are the only prompts known by the dataset owner), the model fine-tuned by original captions can trigger the memorization and generate images with high similarity scores with ${x}_{dup}$ as expected. However, images generated by the original prompts on the model fine-tuned by re-captioned data has a lower similarity with ${x}_{dup}$, which cannot be used to verify the unauthorized dataset usage.
In summary, by pre-processing, EM can be prevented and fails to protect. To overcome the challenges, instead of duplicating data for EM, we propose to use TM to protect the copyright. With the diverse foreground areas, the similarity between templated examples is much lower than the de-duplication threshold, as detailed in Sec. 4.2. Meanwhile, by adjusting the foregrounds, we can make the re-generated captions to have a few shared tokens, which is also able to trigger TM.
@@ 4.1 Template Memorization
    Robustness. The protection should be robust to dataset preprocessing, such as image corruption, noise purification [33] and re-captioning. Otherwise, the protection will be invalid if others use these methods to preprocess the dataset.
@@
    Fast injection. Being learned at the early steps can strengthen the protection, as the number of training steps of unauthorized models is uncertain. Otherwise, if the fine-tuning steps are not enough, TM cannot be injected.
@@
Based on the expectations, in the following subsections, we introduce our framework, EnTruth, from two perspectives, i.e., the generation of template and foregrounds.
@@ 4.2 Generation of Template
  • Step 1: Generating the candidate templates. We utilize SD to generate the candidate templates. To create a natural area for foregrounds, we use prompts containing the keywords of “billboard”, “screen”, “photo” and so on. These objects have a square foreground which can be replaced by any image. The prompts for template can be found in Appd. B.1.
@@ 4.3 Generation of Foregrounds
In summary, based on aforementioned strategies on foregrounds, we can further improve the memorization speed, and the robustness under re-captioning. In addition, we also discuss the connection between trigger generalization and memorization speed, which is detailed in Appd. C.
@@ 4.4 Two Levels of Verification
One-query test involves querying the model only one time and using the classification result to determine whether the model is trained on our dataset. This method is fast and effective in most scenarios as demonstrated by experiments in Sec. 5. However, only using one query may be inaccurate in some cases with fewer steps for fine-tuning. Thus, to get a stable result, we introduce multiple-query test. We can query the model $N(N>1)$ times and use the statistical hypothesis testing in [34, 12] to determine whether the multiple results are significant. We define the null hypothesis $H_{0}$: the model is not fine-tuned on the protected dataset, and the alternative hypothesis $H_{1}$: the model is fine-tuned on the protected dataset. Following [34], we can reject $H_{0}$ at a significant level $\alpha$ if
@@
where $P$ is the number of queries classified as templated in the $N$ queries, $\beta$ is the expected possibility that a non-templated image is wrongly classified by the classifier, $\tau$ is the additional uncertainty margin, and $T_{1-\alpha}$ is the $(1-\alpha)$-quantile of $t$-distribution with $N-1$ degrees of freedom. Different from [34, 12], we use the error rate of the classifier on generated images to estimate $\tau$.
@@ 5 Experiment
Datasets and unauthorized T2I models. We conduct experiments on three datasets, including CC-20k sampled from Conceptual, Captions [7], Sketchyscence [35] with 7265 sketchy images with no caption and Cartoon-blip-caption [36] with 3121 cartoon images captioned by BLIP [32]. We also use BLIP to caption Sketchyscence. More details are in Appd. A.1. We use SD v1.4 and SD v2 as the unauthorized T2I models. Unless otherwise stated, we fine-tune the UNet part of SD for 20,000 steps. We also test with Lora [37] and an online fine-tuning API from OctoAI (https://octo.ai/).
Baselines and metrics. For one-query test, we compare our method with multiple watermark methods, DIAGNOSIS [12], and FT-Shield [11]; poison-only backdoor by dirty label (DL-Backdoor) adapted from [16, 17]. For multiple-query test, we compare the black-box MI by [14]. The details of baselines is in Appd. A.2. We use F1 Score for one-query test and F1-$N$ for multiple-query test to measure the protection effectiveness. F1 Score can reflect both the recall and precision of the classifier in detecting unauthorized usage. F1-$N$ is the F1 Score of detection by multiple-query test with $N=30$ and $\alpha=0.05$. We use FID [38] (calculated on 10,000 generated images) to measure the generation quality.
Implementation details. We use SD to generate templates for CC-20k. For Sketchyscence and Cartoon-blip-caption, since they are not in realistic style, we use an SD fine-tuned on them to generate a template in the sketchy and cartoon domain. Without otherwise stated, we use data-alteration rate of 0.5% for EnTruth, 20% for DIAGNOSIS, 100% for FT-Shield, and 1% for DL-Backdoor. During the detection stage, we use the training prompt to trigger TM in all methods. All the experiments are conducted on a single A5000 GPU.

### v2

# The First Workshop on Generative and Protective AI for Content Creation

@@ 1 Introduction
The latest advancements in generative diffusion models (GDMs) ho2020denoising; song2020denoising; song2020score, especially the text-to-image (T2I) models ramesh2022hierarchical; rombach2022high which excel in creating high-quality images that closely align with the given textual prompts, have revolutionized the field of image generation. These advantages stem not only from the development of model architectures and computing power, but also from the availability of large-scale datasets schuhmann2022laion; sharma2018conceptual; lin2014microsoft. While datasets play an important role, their copyright protection has remained as an unsolved issue. The protection of these datasets’ copyrights is paramount for multiple reasons li2025towards; li2025towards2. For instance, open-source datasets deng2009imagenet are generally available only for educational and research purposes, barring any commercial use. Additionally, for commercial datasets, it is crucial for companies to secure them from theft and unauthorized sales. While pre-training and fine-tuning both raise concerns of copyright infringement, fine-tuning has a more severe impact on the copyright of datasets. Compared to pre-training, fine-tuning is highly efficient, allowing for many unauthorized uses without effective regulatory restrictions.
Observing the above, techniques like watermarking ma2023generative; cui2023ft; wang2023diagnosis; cui2023diffusionshield and black-box Membership Inference (MI) pang2023black; duan2023diffusion have been employed to protect data specifically against unauthorized fine-tuning in text-to-image diffusion models. Nevertheless, existing watermark methods often face some common problems. For example, they usually modify a large portion wang2023diagnosis or even the whole of the dataset cui2023ft, which is not realistic for large-scale datasets. They also unexpectedly affect the quality of generation and are not robust enough under image corruption cui2023diffusionshield; cui2023ft. Meanwhile, as black-box MI does not alter the data to boost the detection, it needs highly extensive queries to get a significant result. Another line of techniques, poison-only backdoor attack shan2023prompt; pan2023trojan, can be adapted for detecting dataset usage by verifying the attacked behavior. However, they are inherently designed for malicious attacking and demonstrate reduced robustness when subjected to re-captioning (as shown by Sec 4.2).
@@
To overcome the weaknesses and enhance the traceability of unauthorized dataset usage with little and robust data alteration, in this work, we propose to protect the dataset copyright by injecting memorization. In T2I models, memorization refers to the phenomenon where the models memorize and reproduce training examples when queried by a memorized prompt carlini2023extracting; naseh2023memory; somepalli2023understanding. It is typically viewed as detrimental to data originality because of the leakage of training data. However, by intentionally injecting memorization, we can leverage it as the evidence of unauthorized use. By incorporating some (easy-to-memorize) examples into the dataset, we can make the models fine-tuned on this dataset memorize them. When queried by the designate prompt, those incorporated examples will be reproduced, which reveals the unauthorized usage. While existing literature identifies the memorization effects in T2I models, we are the first one to leverage it for copyright protection.
@@
According to whether the training examples are partially or entirely memorized, memorization can be divided into exact memorization (EM) and template memorization (TM) webster2023reproducible; ren2024unveiling. To compare EM and TM, EM is the easier one to inject since it is found that simple duplicate data can cause EM carlini2023extracting; wen2023detecting. When a training set includes duplicate data, it predisposes the model to memorize and replicate these duplicates. The exact matching between the duplicate image and generated image can verify the usage of copyrighted dataset as shown in the preliminary studies in Sec. 2. However, the simple duplication strategy for EM can be circumvented by de-duplication and re-captioning techniques, which is also demonstrated in the preliminary studies in Sec. 2. In terms of TM, as shown in Fig. 1, the memorized training images share a common region (named as template), while their remaining areas (named as foreground) differ. Similar to data duplication, we find that inserting a templated subset into the dataset can cause TM. Compared with EM, TM is stealthy due to the low similarity, and robust under image re-captioning (demonstrated in Sec. 3.2 and Sec. 4.2).
@@
Observing the above difference between EM and TM, to generate a stealthy and effective templated set, we propose a novel framework, EnTruth, which Enhances the Traceability of unauthorized dataset usage by TM. Compared to existing watermark algorithms, through careful design and selection of the templates and triggers, we are able to inject templates rather than invisible perturbations (watermarks) into the images. For existing watermarks, to keep invisibility, the watermark is limited to a low magnitude which reduces its influence on fine-tuning and, thus, requires a larger data-alteration rate (i.e. modifying more data samples) as compensation. Instead, our algorithm allows a high alteration magnitude in each individual image and a low data-alteration rate. With such a design, we also enjoy two benefits. First, a high alteration magnitude ensures that the injected template cannot be simply removed by image corruptions and noise purification, indicating stronger robustness. Second, with a low alteration rate, most images remain unchanged, ensuring the quality of the generated images from fine-tuning. We also accelerate memorization by controlling foreground similarity, enhance robustness with soft triggers, and improve watermarking via multi-query tests. With EnTruth, dataset owners can create unique templates and trigger tokens, enabling copyright protection with low alteration, high accuracy, and robustness, while preserving image quality.
@@ 2 Preliminary Study
As mentioned in Section 1, memorization is a common phenomenon in GDMs, and we propose to leverage it in dataset protection. Depending on whether the generative images are totally or partially matching with the training images, memorization can be categorized into exact memorization (EM) and template memorization (TM), and the causes of them are different ren2024unveiling. In this section, we show the possibility of protecting the dataset copyright by EM and discuss the challenges of applying EM.
@@ 2.1 Exact memorization by data duplication enhances the detection of unauthorized usage
Data duplication has been found as one important cause for exact memorization carlini2023extracting; somepalli2023understanding. By duplicating a specific data sample in the training set, the model can accurately memorize and generate it wen2023detecting; ren2024unveiling. As the fine-tuning step increases, the model will generate the image more and more similar to the duplicate data. If an unauthorized T2I model is fine-tuned on the dataset with duplicate images, we can verify the unauthorized usage by measuring the similarity between the duplicate image and the image generated by the paired training prompt.
In Fig. 2(a), we demonstrate the change of similarity score (measured by SSCD pizzi2022self) of duplicate data. We fine-tune Stable Diffusion (SD) starting from the checkpoint v1.4 using CC-20k, a subset of 20,000 text-image pairs from Conceptual Captions sharma2018conceptual. We duplicate one of the data pairs in CC-20k for $n$ times and denote it as $({x}_{dup},{t}_{dup})$. Usually, a larger $n$ can cause memorization with fewer steps. In Fig. 2(a), we use $n=32$. We denote other non-duplicate data as (x,t)({x},{t}). We compare the similarity score between training images and images generated by ${t}_{dup}$ and t{t}. In Fig. 2(a), the similarity score of duplicate data increases much faster than non-duplicate data. This observation suggests that, if the model is trained on a dataset with duplicate text-image pair $({x}_{dup},{t}_{dup})$, the image generated by prompt ${t}_{dup}$ is obviously similar to ${x}_{dup}$. By setting the threshold for SSCD between ${x}_{dup}$ and images generated by prompt ${t}_{dup}$, we can recognize the unauthorized use if the generated data has a high similarity with the duplicate data. Consequently, EM can achieve an accuracy of 74.5% at 10,000 fine-tuning steps with a threshold of 0.1 and 100% at 20,000 steps with a threshold of 0.2.
@@ 2.2 Challenges of Data Duplication
Data de-duplication. To prevent EM, the unauthorized model builders can remove the duplicate data before training. For example, Somepall et al. somepalli2023understanding calculate the similarity score, SSCD pizzi2022self, of each pair of training images, and remove the cluster connected by high similarity scores. In Fig. 2(b), we plot SSCD of natural non-duplicate images. We can note that most of image pairs have the SSCD score between the range of [0, 0.2], while the duplicate data samples have the SSCD of 1. By setting a threshold of 0.7, which is a threshold commonly used to recognize identical images somepalli2023understanding; webster2023reproducible; wen2023detecting, all the duplicate data can be easily removed and no EM can be detected in generated images. Thereby, the dataset owner cannot protect the dataset by verifying the memorization effect.
Image re-captioning. EM relies on the memorized prompts to trigger the memorization. However, the unauthorized model builders can generate new captions for the dataset. Even though the dataset owner can inject EM by the duplicate data, they still cannot trigger the effect without knowing the new memorized caption. We generate new captions for cc-20k by BLIP li2022blip, and fine-tune SD using the original dataset and the re-captioned dataset, respectively. In Fig. 2(c), we calculate SSCD between generated images and ${x}_{dup}$. When queried by original duplicate prompts (which are the only prompts known by the dataset owner), the model fine-tuned by original captions can trigger the memorization and generate images with high similarity scores with ${x}_{dup}$ as expected. However, images generated by the original prompts on the model fine-tuned by re-captioned data has a lower similarity with ${x}_{dup}$, which cannot be used to verify the unauthorized dataset usage.
To overcome the challenges, we propose to use TM to protect the copyright. With the diverse foreground areas, the similarity between templated examples is much lower than the de-duplication threshold, as detailed in Sec. 3.2. Meanwhile, by adjusting the foregrounds, we can make the re-generated captions to have a few shared tokens, which is also able to trigger TM.
@@ 3.1 Template Memorization
    Robustness. The protection should be robust to dataset preprocessing, such as image corruption, noise purification naseer2020self and re-captioning. Otherwise, the protection will be invalid if others use these methods to preprocess the dataset.
@@
    Fast injection. Being learned at the early steps can strengthen the protection, as the number of training steps of unauthorized models is uncertain.
@@ 3.2 Generation of Template
  • Step 1: Generating the candidate templates. We utilize SD to generate the candidate templates. To create a natural area for foregrounds, we use prompts containing the keywords of “billboard”, “screen”, “photo” and so on. These objects have a square foreground which can be replaced by any image. The prompts for template can be found in Appd. C.1.
@@ 3.3 Generation of Foregrounds
In summary, based on aforementioned strategies on foregrounds, we can further improve the memorization speed, and the robustness under re-captioning. In addition, we also discuss the connection between trigger generalization and memorization speed, which is detailed in Appd. D.
@@ 3.4 Two Levels of Verification
One-query test involves querying the model only one time and using the classification result to determine whether the model is trained on our dataset. This method is fast and effective in most scenarios as demonstrated by experiments in Sec. 4. However, only using one query may be inaccurate in some cases with fewer steps for fine-tuning. Thus, to get a stable result, we introduce multiple-query test. We can query the model $N(N>1)$ times and use the statistical hypothesis testing in li2023black; wang2023diagnosis to determine whether the multiple results are significant. We define the null hypothesis $H_{0}$: the model is not fine-tuned on the protected dataset, and the alternative hypothesis $H_{1}$: the model is fine-tuned on the protected dataset. Following li2023black, we can reject $H_{0}$ at a significant level $\alpha$ if
@@
where $P$ is the number of queries classified as templated in the $N$ queries, $\beta$ is the expected possibility that a non-templated image is wrongly classified by the classifier, $\tau$ is the additional uncertainty margin, and $T_{1-\alpha}$ is the $(1-\alpha)$-quantile of $t$-distribution with N1N-1 degrees of freedom. Different from li2023black; wang2023diagnosis, we use the error rate of the classifier on generated images to estimate $\tau$.
@@ 4 Experiment
Datasets and unauthorized T2I models. We conduct experiments on three datasets, including CC-20k sampled from Conceptual, Captions sharma2018conceptual, Sketchyscence zou2018sketchyscene with 7265 sketchy images with no caption and Cartoon-blip-caption Adler with 3121 cartoon images captioned by BLIP li2022blip. We also use BLIP to caption Sketchyscence. More details are in Appd. B.1. We use SD v1.4 and SD v2 as the unauthorized T2I models. Unless otherwise stated, we fine-tune the UNet part of SD for 20,000 steps. We also test with Lora hu2021lora and an online fine-tuning API from OctoAI (https://octo.ai/).
Baselines and metrics. For one-query test, we compare our method with multiple watermark methods, DIAGNOSIS wang2023diagnosis, and FT-Shield cui2023ft; poison-only backdoor by dirty label (DL-Backdoor) adapted from shan2023prompt; pan2023trojan. For multiple-query test, we compare the black-box MI by pang2023black. The details of baselines is in Appd. B.2. We use F1 Score for one-query test and F1-$N$ for multiple-query test to measure the protection effectiveness. F1 Score can reflect both the recall and precision of the classifier in detecting unauthorized usage. F1-$N$ is the F1 Score of detection by multiple-query test with $N=30$ and $\alpha=0.05$. We use FID heusel2017gans (on 10,000 images) to measure the generation quality.
Implementation details. We use SD to generate templates for CC-20k. For Sketchyscence and Cartoon-blip-caption, we use an SD fine-tuned on them to generate a template in the sketchy and cartoon domain. Without otherwise stated, we use data-alteration rate of 0.5% for EnTruth, 20% for DIAGNOSIS, 100% for FT-Shield, and 1% for DL-Backdoor. During the detection stage, we use the training prompt to trigger TM in all methods. All the experiments are conducted on an A5000 GPU.